# MSDiagnosis: An EMR-based Dataset for Clinical Multi-Step Diagnosis

### v1

@@ Appendix B Details of the Evaluation Metrics
For the disease entities in the diagnosis results $D$ and the reference results $R$ in the medical records, we compute the edit distance to associate these entities with ICD-10 terms, thereby mapping $D$ and $R$ to two standardized disease sets, $S_{d}$ and $S_{r}$, respectively. We then compute the entity $F1$ score based on $S_{d}$ and $S_{r}$.
For primary and final diagnosis criteria, we employ two types of evaluation metrics. First, we adapt $Rouge-L$ Lin (2004) and $BLEU$-1 Papineni et al. (2002) to measure the longest matching sequence between the generated text and the reference text, capturing the similarity in the overall structure of the two sequences. Second, we introduce the $Macro-Recall$ metric, which is calculated based on the key information in the answers. For our predefined $N$ scoring point categories, the calculation method for key information $Macro-Recall$ is as follows:
@@ Appendix C Details of Implementation
To enhance the stability and reliability of the experimental results and reduce the impact of random factors, we conduct each experiment three times and then calculate the average of three results. For the backbone model, we utilize the OpenAI API, specifying the model as “GPT4o-mini”. We set the $top\_p$ parameter to 0.01, and all other hyperparameters of the OpenAI API are maintained at default values. When selecting similar examples, we set $K$ to 1.
@@ Appendix D Question Definition
Table 7: The definition of question
In this section, we introduce in detail the definition of the diagnostic questions corresponding to each EMR. In ReaDiagnosis, the questions corresponding to each case are initially constructed manually and then expanded using GPT-4. The definitions of each question are shown in Table 7.
Number of Disease F1 Macro-Recall Rouge-L Blue
Pre DD Fin Pre Fin Pre Fin Pre Fin
1-5 35.93 11.63 32.50 42.52 45.73 57.33 58.4 50.17 48.62
5-10 30.03 9.77 26.46 31.73 22.47 49.76 45.92 45.39 45.16
>10 34.27 9.35 29.18 17.94 22.57 48.82 50.19 45.95 42.86
Table 8: The impact of the number of diagnosed diseases on the framework.
@@ Appendix E The Impact of Different Departments on the Framework
In this section, to evaluate the impact of different departments on our framework, we conduct an analysis of the department-specific results. The detailed experimental outcomes are presented in Table 6. From this analysis, we can draw the following conclusions: Different departments have different influences on the framework. The Oncology and Traditional Chinese Medicine departments have a significant impact on our framework. Specifically, the final $F1$ score in the Oncology department is only 5.89%.
@@ Appendix G The Impact of the Number of ICL on the Framework
Model F1 Macro-Recall Rouge-L Blue
Pre DD Fin Pre Fin Pre Fin Pre Fin
ICL=0 34.70 10.09 31.91 34.58 31.91 54.25 53.31 47.04 44.66
ICL=1 34.78 10.97 31.69 40.67 42.28 55.95 55.97 49.15 47.94
ICL=2 34.76 10.72 31.53 41.60 43.88 56.92 55.28 50.42 48.72
ICL=3 35.37 10.73 31.69 42.45 43.08 54.16 54.94 50.49 49.31
Appendix H Error Case
In this section, we present an error example in Fig. 4 to facilitate a better understanding of the error categories. The details are as follows: 1) Lack of domain knowledge (42%): For example, based on the patient’s medical history and examination results, the model fails to accurately predict the patient’s disease. This error is likely due to the model’s limited domain knowledge, which leads to the model not accurately predicting the disease. 2) Symptoms or examination results are confused with the disease (30%): For example, “Intracranial Space-Occupying Lesion” and “Tinnitus” are the patient’s examination results and symptoms, respectively. However, the model predicts them as diseases. 3) Diagnostic basis inconsistent with facts (18%): For example, the model incorrectly extracts the patient’s blood pressure as 140/80.
Figure 4: Case study. The green (red) highlight indicates correct (incorrect) results.
Appendix I The Prompt used in our Framework
In this section, we primarily introduce the prompts corresponding to the four instances where the LLM is used within the framework. The prompt for forward reasoning is shown in Fig. 5. The prompt for backward reasoning is shown in Fig. 7. The prompt for reflecting on the diagnostic results is shown in Fig. 8. The prompt for optimizing the diagnostic results is shown in Fig. 6.
@@ Appendix I The Prompt used in our Framework
Figure 8: The prompt of reflection on diagnostic result.

### v3

@@ Appendix B Details of the Evaluation Metrics
For the disease entities in the diagnosis results $D$ and the reference results $R$ in the medical records, we employ a more rigorous evaluation method. This approach aligns better with actual clinical needs and facilitates targeted treatment for patients. Specifically, we compute the edit distance to associate these entities with ICD-10 terms, thereby mapping $D$ and $R$ to two standardized disease sets, $S_{d}$ and $S_{r}$, respectively. We then compute the entity $F1$ score based on $S_{d}$ and $S_{r}$.
For primary and final diagnosis criteria, we employ two types of evaluation metrics. First, we adapt $Rouge-L$ Lin (2004), $BERTScore$ Zhang* et al. (2020), and $BLEU$-1 Papineni et al. (2002) to measure the longest matching sequence between the generated text and the reference text, capturing the similarity in the overall structure of the two sequences. Second, we introduce the $Macro-Recall$ metric, which is calculated based on the key points in the answers. For our predefined $N$ key point categories, the calculation method for key point $Macro-Recall$ is as follows:
@@ Appendix C Details of Implementation
To enhance the stability and reliability of the experimental results and reduce the impact of random factors, we conduct each experiment three times and then calculate the average of three results. For the backbone model, we utilize the OpenAI API, specifying the model as “GPT4o-mini”. We set the $top\_p$ parameter to 0.01, and all other hyperparameters of the OpenAI API are maintained at default values. When selecting similar examples, we set $K$ to 1. During the LoRA fine-tuning phase, we employ the Adam optimizer with weight decay correction for fine-tuning the model. The initial learning rate is set to 1e-4, and the batch size is set to 1, with Cross-Entropy Loss serving as the loss function. The input during training is: "instruction: task description. Medical Record: patient case. question: diagnosis question". The output is: "the answer to the question". We calculate the loss only for the answer, excluding the instruction and input. For the open-source models, our experiments are conducted on four Nvidia A100 GPUs, each with 40GB of memory, and we use PyTorch [^3] https://pytorch.org/ in Python [^4] https://www.python.org/.
@@ Appendix D Question Definition
In this section, we introduce in detail the definition of the diagnostic questions corresponding to each EMR. In MSDiagnosis, the questions corresponding to each case are initially constructed manually and then expanded using GPT-4. The definitions of each question are shown in Table 7.
Table 8: The impact of the number of diagnosed diseases (final diagnosis) on the framework.
Number of Disease F1 Macro-Recall Rouge-L BLEU-1 BERTScore
Pre DD Fin Pre Fin Pre Fin Pre Fin Pre Fin
1-5 35.93 11.63 32.50 42.52 45.73 57.33 58.4 50.17 48.62 86.32 86.23
5-10 30.03 9.77 26.46 31.73 22.47 49.76 45.92 45.39 45.16 84.58 84.75
>10 34.27 9.35 29.18 17.94 22.57 48.82 50.19 45.95 42.86 84.66 86.17
@@ Appendix E The Impact of Different Departments on the Framework
In this section, to evaluate the impact of different departments on our framework, we conduct an analysis of the department-specific results. The detailed experimental outcomes are presented in Table 6. From this analysis, we can draw the following conclusions: Different departments have different influences on the framework. The Oncology have the greatest impact on our framework. Specifically, the final $F1$ score in the Oncology department is only 5.89%. The best performance is achieved in the Stomatology department. Specifically, the final $F1$ score in the Oncology department is only 83.33%.
@@ Appendix F The Impact of the Number of Diagnosed Diseases on the Framework
Model F1 Macro-Recall Rouge-L BLEU-1 BERTScore
Pre DD Fin Pre Fin Pre Fin Pre Fin Pre Fin
ICL=0 34.70 10.09 31.91 34.58 31.91 54.25 53.31 47.04 44.66 84.30 84.45
ICL=1 34.78 10.97 31.69 40.67 42.28 55.95 55.97 49.15 47.94 85.19 85.72
ICL=2 34.76 10.72 31.53 41.60 43.88 56.92 55.28 50.42 48.72 86.45 86.04
ICL=3 35.37 10.73 31.69 42.45 43.08 54.16 54.94 50.49 49.31 87.19 87.72
@@ Appendix G The Impact of the Number of ICL on the Framework
Figure 7: The prompt of reflection on diagnostic result.
Appendix H The Impact of Different Parameters of the Model on the Framework
To evaluate the impact of models with varying parameter scales on the framework’s performance, we evaluate Qwen2.5 models with 7B, 14B, 32B, and 72B parameters, and also compare with the 1-shot method for analysis convenience. The detailed experimental results are presented in Table 10. From the table, the following conclusions emerge: as the number of parameters increases, the framework’s performance improves progressively. Specifically, the 32B model outperforms the 7B model in the F1 and Macro-Recall metrics for final diagnosis by 2.18% and 6.96%, respectively. In comparison to the 14B model, the 32B model achieves improvements of 1.26% and 1.12% in these metrics.
Table 10: Comparison of models with different parameter sizes on our framework.
Model Method F1 Macro-Recall Rouge-L BLEU-1 BERTScore
Pre DD Fin Pre Fin Pre Fin Pre Fin Pre Fin
Qwen2.5-7B-Instruct 1-Shot 34.65 10.56 29.97 34.90 36.07 53.19 54.41 41.73 42.26 86.67 87.21
Our 36.97 9.01 33.35 39.39 40.23 58.60 59.81 50.32 49.25 87.61 87.57
Qwen2.5-14B-Instruct 1-Shot 36.88 11.72 33.82 43.12 43.18 60.27 60.61 53.06 50.48 88.68 89.84
Our 36.95 12.41 34.27 44.31 46.07 63.07 61.04 53.17 52.01 89.03 89.56
Qwen2.5-32B-Instruct 1-Shot 37.43 11.92 33.40 43.46 45.82 52.99 51.16 38.07 36.45 85.72 85.61
Our 38.27 11.79 35.53 45.24 47.19 62.83 63.38 55.43 52.64 88.37 88.48
Qwen2.5-72B-Instruct 1-Shot 36.87 10.56 34.11 47.41 50.01 59.34 60.11 52.51 52.24 87.27 87.83
Our 38.76 10.56 34.97 49.24 51.09 64.71 65.10 54.14 53.64 87.52 88.83
Appendix I The Prompt used in our Framework
In this section, we primarily introduce the prompts corresponding to the four instances where the LLM is used within the framework. It is noteworthy that the prompts in our framework follow a general pattern, including role definition, formatting instructions, and examples, rather than being meticulously designed. The prompt for forward reasoning is shown in Fig. 5. The prompt for backward reasoning is shown in Fig. 8. The prompt for reflecting on the diagnostic results is shown in Fig. 7. The prompt for optimizing the diagnostic results is shown in Fig. 6.
@@ Appendix I The Prompt used in our Framework
Appendix J Complete Multi-step Diagnostic Example
We present a comprehensive multi-step diagnostic example, as illustrated in Fig. 9. In the first stage, the input is the medical record introduction, and Q1 and Q2 are answered sequentially. The responses to these questions are A1 and A2. In the second stage, the input includes the medical record introduction and the Q&A history from the first stage, and Q2 must be addressed. In the third stage, the input consists of the medical record introduction, the Q&A history from the first two stages, and the patient’s diagnosis and treatment process. In this stage, Q4 and Q5 must be answered sequentially.
Figure 9: A complete multi-step diagnostic example.